OpenStudy (anonymous):

Solve 4 + 12a = -2(6-4a)

OpenStudy (saifoo.khan):

4 + 12a = - 12a + 8a

OpenStudy (anonymous):

This time you have to first distribute: 4 + 12a = -2(6-4a) 4 + 12a = -12 + 8a Then subtract 8a 4 + 4a = -12 subtract 4 4a = -16 divide by 4 a = -4